OpenStudy (anonymous):

Tell whether the sequence 1/3,0 ,1 ,-2 . . . is arithmetic, geometric, or neither. Find the next three terms of the sequence. A. neither; 7, -20, 61 B. geome

OpenStudy (anonymous):

for a geometric series the common ration should be the same for two terms like r=n2/n1=n3/n2 r=0/(1/3) is not equal to 1/0 so it is not a geometrc series

OpenStudy (anonymous):

for a arithmetic series the common difference ashould be the same d=n2-n1=n3-n2 d=0-1/3 is not equal to 1-0 conclusion it is neither arithametic nor geometric

OpenStudy (anonymous):

I guess the only option is A. for neither. The 7, 20, 61 look arbitrary.

OpenStudy (anonymous):

Thats what i thought it was but wasent sure thank you

OpenStudy (anonymous):

Np ;) Just hit on the blue button next to me. It helps me out alot. :b